OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

Here's a different way: (sort of) The slope of the line is \[(y-y_1)/(x-x_1)\], so slope=9/3=3 10=3*1+b b=7 y=3x+7 which is a different way of writing 3x-y=-7

OpenStudy (anonymous):

alright thanks guys!